OpenStudy (alesia/bailey):

what's the point slope formula, Slope-intercept formula and standard form formula??

OpenStudy (anonymous):

point slope y – y1 = m(x – x1) slope intercept is y=mx+b and standard is ax+by=c

OpenStudy (anonymous):

refresh the page to see it if it doesnt come out right

OpenStudy (alesia/bailey):

okay. thanks ^^